a 100 lb bag of a rye/blue grass mix sells for $245. a 100 lb bag of blue grass sells for $347. if 19 bags combined sells for $5369 how many of each are sold
Answer by ptaylor(2198)

SOLVE USING 2 EQUATIONS AND TWO UNKNOWNS:
Let x=number of bags of rye/blue sold
And y=number of bags of blue grass sold
Now we are told the following:
x+y=19------------------------------eq1
$ understood
245x+347y=5369--------------------eq2
from eq1, x=19-y substitute 19-y for x in eq2:
245(19-y)+347y=5369 get rid of parens:
4655-245y+347y=5369 subtract 4655 from each side
4655-4655-245y+347y=5369-4655 collect like terms
102y=714 divide each side by 102
y=7---------------------number of bags of blue grass sold
substitute y=7 into eq1
x+7=19 subtract 7 from each side
x=12--------------------------number of bags of blue/rye sold
SOLVE USING 1 EQUATION AND 1 UNKNOWN:
Let x=number of bags of blue/rye
Then 19-x=number of bags of blue
So:
245x+347(19-x)=5329
And this will yield the same answer as before
Hope this helps----ptaylor